# Control the GNN: Utilizing Neural Controller with Lyapunov Stability for Test-Time Feature Reconstruction

### v1

@@ Abstract
The performance of graph neural networks (GNNs) is susceptible to discrepancies between training and testing sample distributions. Prior studies have attempted to enhance GNN performance by reconstructing node features during the testing phase without modifying the model parameters. However, these approaches lack theoretical analysis of the proximity between predictions and ground truth at test time. In this paper, we propose a novel node feature reconstruction method grounded in Lyapunov stability theory. Specifically, we model the GNN as a control system during the testing phase, considering node features as control variables. A neural controller that adheres to the Lyapunov stability criterion is then employed to reconstruct these node features, ensuring that the predictions progressively approach the ground truth at test time. We validate the effectiveness of our approach through extensive experiments across multiple datasets, demonstrating significant performance improvements.
Introduction
@@
Graph neural networks (GNNs) have gained significant attention in recent years due to their strong representational capabilities (Wu et al. 2020; Yuan et al. 2022). As the applications and data scales of GNNs continue to grow, their performance becomes sensitive to discrepancies between training and testing sample distributions(Qin et al. 2022; Li et al. 2022). To address performance degradation caused by distribution shift, some approaches enhance the model’s generalization during the training phase using techniques such as data augmentation(Sui et al. 2024; Ding et al. 2022), causal inference(Fan et al. 2023), and regularization(Buffelli, Liò, and Vandin 2022). However, these techniques require modifications to the model structure or parameters, limiting their applicability when the model architecture is fixed or when retraining is prohibitively expensive.
To improve the practicality of graph neural networks, some approaches focus on reconstructing node features during the testing phase to enhance model performance without altering parameters. Well-designed perturbations to GNN input data  (Jin et al. 2023) is introduced to reduce the effects of distribution shift, guided by a surrogate loss. However, this method results in node features with low feasibility. To address this, FRGNN  (Ding et al. 2024) employs an MLP to learn the mapping between GNN outputs and inputs, generating class representative embeddings for each node class during the testing phase. Then FRGNN replaces the features of known label nodes with these class representative embeddings to enhance model robustness against distribution shift. Despite their effectiveness, this method relies on manually designed rules or mechanisms for node feature reconstruction and lacks theoretical analysis of the proximity between predictions and ground truth at test time.
To improve the feasibility of feature reconstruction methods during the testing phase, we formalize GNNs as control systems, treating node features as controlled variables and the predictions as state. The aforementioned methods can be viewed as solving an optimization problem, specifically finding a node feature that minimizes the discrepancy between the GNN’s prediction and the ground truth. Transforming this problem into a control problem offers the following advantages:
  • Compared to the approaches of optimizing a single loss function, this method offers more precise control over each individual state.
  • Introducing a control system allows the use of stability criteria from control theory to analyze the stability of the system, determining whether the system output remains stable near the equilibrium point.
Our goal is to find a controller to ensure the predictions get close to the ground truth gradually. However, due to the complex nonlinearity of GNNs, traditional controllers face significant challenges in meeting the control requirements. Methods like the proportional–integral–derivative controller (PID)  (Ang, Chong, and Li 2005) and the linear quadratic regulator (LQR)  (Kwakernaak and Sivan 1972) are widely used in various fields. However, the control effectiveness of these methods for nonlinear systems is relatively poor. In contrast, model-free adaptive control (MFAC)  (Hou and Xiong 2019) has superior capability in handling nonlinear dynamics. But unlike ordinary neural networks like MLPs, GNNs include a message-passing layer. Adjusting the features of specific nodes affects not only the output of those nodes but also influences the prediction outputs of other nodes through the graph structure. This interdependence makes the control problem more complex and challenging for such controllers. To tackle this issue, we introduce a neural controller. Neural controllers have been successfully used in various tasks, including inverted pendulum and drone control  (Jiang, Pourpanah, and Hao 2020; Kim and Calise 1994; Gu et al. 2020). To ensure stability, some studies have proposed neural Lyapunov functions for neural controllers  (Chang, Roohi, and Gao 2020; Zhou et al. 2022; Dai et al. 2021). However, these approaches are focused on traditional control problems and have not been applied to the node feature reconstruction problem in GNNs.
To broaden the application scope of neural controllers, we apply them to the node feature reconstruction in graph neural networks. We utilize neural controllers to adjust node features and learn controller parameters in a data-driven manner. This method aims to control the node’s predicted output to the ground truth gradually. To ensure that the node prediction output remains stable near the ground truth, we require the learned neural controller to satisfy the Lyapunov stability criterion. Since it is challenging to design Lyapunov functions for complex controllers manually, we introduce a neural Lyapunov function to approximate the Lyapunov function through a neural network. We simultaneously learn the parameters of the neural controller and the neural Lyapunov function to ensure that the controller satisfies the Lyapunov criterion under this Lyapunov candidate, thereby ensuring the stability of the control system. Our contributions are summarized as follows:
  • We propose a new paradigm for applying control theory to enhance the performance of graph neural networks. We formalize the GNN as a control system, considering node features as controlled variables and the predictions as state. We design controllers to optimize these variables to improve model prediction results.
  • We design a neural controller suitable for the performance enhancement tasks of graph neural networks. To ensure that the controller can effectively adjust node features, we introduce a neural Lyapunov function. By ensuring the validity of this Lyapunov function, we guarantee the stability of the controller.
  • We validate the effectiveness of our method on multiple datasets. Experimental results demonstrate that our approach significantly enhances the performance of graph neural networks by mitigating the impact of distribution shift, all without altering the model parameters.
Related Work
Test-Time Feature Reconstruction For GNNs
@@
In the task of semi-supervised node classification, GNNs often suffer from a distribution shift between the training set and testing set due to the inappropriate selection of training samples or limited training data. To tackle this issue without altering the model structure and parameters, reconstructing node features during the testing phase is effective. GTrans (Jin et al. 2023) employs surrogate loss to guide the update of perturbations during the testing phase to reconstruct node features. FRGNN (Ding et al. 2024) proves that replacing the node features of labeled nodes with class representative embeddings during the testing phase enhances the model’s robustness to distribution shift. Identifying appropriate class representative embeddings is crucial for this method. FRGNN maps the relationship between GNN outputs and inputs using an MLP. Then, it uses the one-hot embedding of the class as input to the MLP to obtain the class representative embeddings. However, the class representative embeddings obtained in this way do not guarantee that the new output of GNN will be close to the one-hot embedding of the corresponding class, which fails to meet the requirements of class representative embeddings. To address this issue, we design a neural controller to find appropriate class representative embeddings by ensuring that the new prediction of GNN gets close enough to the ground truth.
Figure 1: The overall process of our method.
Neural Controller and Neural Lyapunov Function
@@
Lyapunov stability is commonly used in control theory to ensure system stability. However, designing Lyapunov functions of neural controllers is challenging. To address this issue,  (Chang, Roohi, and Gao 2020) presents a learning framework to learn control policies and their corresponding Lyapunov functions. This framework guides model training by continuously searching for counterexamples that do not satisfy the Lyapunov condition until all states in the domain satisfy the Lyapunov condition. However, this method requires the Lyapunov function to be differentiable for counterexample search with the SMT solvers. To meet this requirement, all activation functions in  (Chang, Roohi, and Gao 2020) are Tanh. To relax the restriction that the activation function must be continuous,  (Dai et al. 2021) proposes using Mix Integer Programs (MIP) to find counterexamples. This method allows the widely used ReLU activation function to be used. The authors rigorously prove the validity of the Lyapunov function.  (Zhou et al. 2022) learns the unknown nonlinear system through a neural network without the need for precise system modeling. Then, the system is controlled using the framework proposed in  (Chang, Roohi, and Gao 2020). The authors theoretically guarantee the validity of the Lyapunov function under the unknown system. Although these methods simplify the design of Lyapunov functions, they solve traditional control problems rather than enhancing the neural networks’ performance with Lyapunov functions. Our work is inspired by these studies. We design a neural controller to reconstruct node features to improve model prediction.
Methodology
@@
In this section, we introduce the method of reconstructing node features using a neural controller. The overall process of this method is shown in Figure 1. Initially, we use a well-trained GNN to obtain the node predictions. This output is then fed back into the neural controller to reconstruct the node features, ensuring the new predictions remain close to the ground truth. To ensure that the new predictions derived from the reconstructed node features remain stable near the ground truth, we introduce a neural Lyapunov function to aid in controller training. This section is organized as follows. We first explain the theoretical foundation of this method. Then, we detail the design of the neural controller and the neural Lyapunov function. Finally, we outline the training approach for this neural controller and the neural Lyapunov function.
Problem Formulation
@@
In FRGNN  (Ding et al. 2024), the authors demonstrate that replacing the features of labeled nodes with class representative embeddings during the test phase can improve the model’s robustness to distribution shift. This paper provides the definition of class representative embeddings, as shown in Definition 1.
Definition 1
@@
where $f$ represents a function. Our goal is to identify a function $f$ to minimize the objective function in Equation 3. However, solving this optimization problem is challenging due to the complex nonlinearity of GNNs. We reanalyze the problem. Considering the advantages of converting the optimization problem into a control problem mentioned in our introduction, we transform the optimization challenge into a control problem. We can formalize the graph neural network as a control system, treating node features as controlled variables and the predictions as states. By designing a controller to adjust these controlled variables, we aim to ensure that the node predictions get close to the ground truth, thereby enhancing the model’s performance. The form of this control system is as follows:
@@
We have now transformed the optimization problem in Equation 3 into designing a controller $f$. We aim to ensure that the controller $f$ satisfies the Lyapunov stability criterion, ensuring that the system output remains stable near the equilibrium point (i.e., ground truth).
Theoretical Foundation
In the previous chapter, we transform the GNNs optimization problem into a control problem. Now, the main challenge is designing a controller $f$ to adjust the node features so that the new predictions derived from these reconstructed features get close to the ground truth. Traditional controllers, such as PID, LQR, and AC, are not well-suited for this system due to the complex nonlinearity of graph neural networks. Moreover, unlike ordinary neural networks like MLPs, GNNs include message-passing layers, meaning that modifying a specific node feature not only affects the prediction of that node but also influences the predictions of other nodes through the graph structure. This coupling makes the control problem particularly challenging for traditional controllers. Therefore, we choose a neural controller, leveraging the powerful representation capabilities of neural networks to learn an effective controller suitable for this system. We denote the neural controller as $f_{\theta}$, where $\theta$ represents the trainable parameters. Thus, Equation 5 can be rewritten in closed-loop form as follows:
@@
Theorem 1
Consider a close-loop system $\hat{Y}_{t+1}=C_{f_{\theta}}(\hat{Y_{t}})$ with an equilibrium point at the $Y$, i.e. $C_{f_{\theta}}(Y)=Y$, where $C_{f_{\theta}}$ is a MLP with ReLU activation function. Suppose there exists a continuously differentiable function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies the following conditions:
@@
Proof 1
Please refer to the Appendix A.
Definition 2
Lyapunov Function: Any continuously differentiable function $V$ maps the state space to a scalar value. Lyapunov Stability Criterion: For a close-loop system $\hat{Y}_{t+1}=C_{f_{\theta}}(\hat{Y_{t}})$ with an equilibrium point at the $Y$, the system is asymptotically stable at the $Y$ if there exists a Lyapunov function $V$ that satisfies the conditions in Equation 7.
@@
Figure 2: The architecture of the neural controller and the neural Lyapunov function.
Training Methodology

### v2

@@ Abstract
The performance of graph neural networks (GNNs) is susceptible to discrepancies between training and testing sample distributions. Prior studies have attempted to mitigating the impact of distribution shift by reconstructing node features during the testing phase without modifying the model parameters. However, these approaches lack theoretical analysis of the proximity between predictions and ground truth at test time. In this paper, we propose a novel node feature reconstruction method grounded in Lyapunov stability theory. Specifically, we model the GNN as a control system during the testing phase, considering node features as control variables. A neural controller that adheres to the Lyapunov stability criterion is then employed to reconstruct these node features, ensuring that the predictions progressively approach the ground truth at test time. We validate the effectiveness of our approach through extensive experiments across multiple datasets, demonstrating significant performance improvements.
Index Terms:
graph neural network, Lyapunov stability, neural controller
I Introduction
@@
Graph neural networks (GNNs) have demonstrated significant success in various applications due to their powerful representational capabilities [1, 2]. However, their performance suffers significantly when distribution discrepancies exist between training and testing samples [3, 4]. Existing approaches often tackle distribution shifts during the training phase through data augmentation [5, 6], causal inference[7], and regularization[8]. Nevertheless, these strategies necessitate modifications to the model structure and relearning the parameters, limiting their applicability when retraining is costly or impossible.
To address this limitation, recent methods propose reconstructing node features during the testing phase, aiming to mitigate distribution shifts without altering the original model parameters. Notably, the term ”node features” here refers explicitly to the original input features of the nodes. For instance, GTrans [9] generates perturbed node features guided by a surrogate loss to minimize distribution shift effects, but these reconstructed features often lack practical feasibility due to arbitrary perturbations. FRGNN [10] addresses this by employing a multi-layer perceptron (MLP) to learn class-representative embeddings, substituting the original node features to enhance robustness. However, both methods rely heavily on heuristic or manually designed mechanisms, lacking theoretical guarantees regarding the proximity of reconstructed features to their ideal representations. Consequently, the interpretability and reliability of these feature reconstruction methods remain limited.
To overcome these limitations, we propose to view the node feature reconstruction problem through the lens of control theory. Specifically, we formalize GNN predictions as system states and node features as controllable inputs. Under this framework, reconstructing node features becomes a control problem, where the goal is to adjust node features precisely to minimize the discrepancy between the GNN predictions and the ground truth. Compared to heuristic optimization approaches, this perspective offers distinct advantages:
  • It enables precise control over individual node states, allowing fine-grained optimization of prediction accuracy.
  • It facilitates stability analysis through established stability criteria from control theory, such as Lyapunov stability, ensuring that feature adjustments do not introduce instability or degrade prediction performance.
However, applying traditional control methods, such as proportional-integral-derivative (PID) controllers [11] or linear quadratic regulators (LQR) [12], faces significant challenges in GNNs. Due to the complex nonlinearity and inherent message-passing mechanisms, adjusting a node’s feature impacts not only its own prediction but also influences other nodes’ predictions through the graph structure. This complex interdependency renders traditional control methods inadequate.
To address this challenge, we employ a data-driven approach using neural controllers, capable of learning complex and nonlinear control policies directly from data. Neural controllers have demonstrated effectiveness in controlling complex nonlinear systems, such as inverted pendulums and unmanned aerial vehicles [13, 14, 15]. However, guaranteeing the stability of neural controllers in highly interconnected systems like GNNs is non-trivial. Manually designing Lyapunov functions for these nonlinear systems is typically difficult or even infeasible. Therefore, we introduce a neural Lyapunov function to simultaneously learn with the neural controller in a data-driven manner. By co-optimizing both the neural controller and the neural Lyapunov function, we ensure that the controller adheres to Lyapunov stability criteria, guaranteeing stable and reliable adjustments of node features during testing.
Although neural Lyapunov methods have shown success in traditional control tasks [16, 17, 18], their application to GNN node feature reconstruction remains unexplored. Our approach extends neural control methods into the domain of graph neural networks, providing a novel solution for theoretically-grounded node feature reconstruction.
  • We present a novel control-theoretic paradigm to enhance GNN robustness against distribution shifts without altering model parameters. By formalizing node predictions as system states and node features as controlled variables, we introduce theoretically stability analysis into test-phase feature reconstruction.
  • We design a neural controller specifically tailored to the nonlinear, interconnected structure of GNNs, capable of effectively reconstructing node features at test time. To guarantee stability during this reconstruction process, we further propose a neural Lyapunov function learned simultaneously with the neural controller, providing theoretical stability assurances for our method.
  • We empirically validate our proposed framework on multiple benchmark datasets. The experimental results demonstrate that our method significantly improves GNN performance under distribution shift scenarios, effectively reconstructing meaningful node features and maintaining prediction stability without modifying the original model structure or retraining its parameters.
II Related Work
II-A Test-Time Feature Reconstruction For GNNs
@@
In the task of semi-supervised node classification, GNNs often suffer from a distribution shift between the training set and testing set due to the inappropriate selection of training samples or limited training data. To tackle this issue without altering the model structure and parameters, reconstructing node features during the testing phase is effective. GTrans [9] employs surrogate loss to guide the update of perturbations during the testing phase to reconstruct node features. FRGNN [10] proves that replacing the node features of labeled nodes with class representative embeddings during the testing phase enhances the model’s robustness to distribution shift. Identifying appropriate class representative embeddings is crucial for this method. FRGNN maps the relationship between GNN outputs and inputs using an MLP. Then, it uses the one-hot embedding of the class as input to the MLP to obtain the class representative embeddings. However, the class representative embeddings obtained in this way do not guarantee that the new output of GNN will be close to the one-hot embedding of the corresponding class, which fails to meet the requirements of class representative embeddings. To address this issue, we design a neural controller to find appropriate class representative embeddings by ensuring that the new prediction of GNN gets close enough to the ground truth.
Figure 1: The overall process of our method. We input the original prediction $\hat{Y}t$ into the neural controller to obtain adjusted node features, and then feed these adjusted node features back into the well-trained GNN to obtain the prediction $\hat{Y}{t+1}$ at time $t+1$. Simultaneously, we input the original prediction $\hat{Y}_{t}$ into the neural Lyapunov function to obtain the Lyapunov value. We then use these values to compute the Lyapunov loss and update the parameters of both the neural controller and the neural Lyapunov function. Afterward, we use a solver to check whether there are states in the state space that do not satisfy the Lyapunov conditions. If such states exist, we add them to the training set and continue training until none remain.
II-B Neural Controller and Neural Lyapunov Function
@@
Lyapunov stability is commonly used in control theory to ensure system stability. However, designing Lyapunov functions of neural controllers is challenging. To address this issue,  [16] presents a learning framework to learn control policies and their corresponding Lyapunov functions. This framework guides model training by continuously searching for counterexamples that do not satisfy the Lyapunov condition until all states in the domain satisfy the Lyapunov condition. However, this method requires the Lyapunov function to be differentiable for counterexample search with the satisfiability modulo theories (SMT) solvers. To meet this requirement, all activation functions in  [16] are Tanh. To relax the restriction that the activation function must be continuous,  [18] proposes using Mix Integer Programs (MIP) to find counterexamples. This method allows the widely used ReLU activation function to be used. The authors rigorously prove the validity of the Lyapunov function.  [17] learns the unknown nonlinear system through a neural network without the need for precise system modeling. Then, the system is controlled using the framework proposed in  [16]. The authors theoretically guarantee the validity of the Lyapunov function under the unknown system. Although these methods simplify the design of Lyapunov functions, they solve traditional control problems rather than enhancing the neural networks’ performance with Lyapunov functions. Our work is inspired by these studies. We design a neural controller to reconstruct node features to improve model prediction.
TABLE I: Summary of variables and their meanings.
Symbol Definition / Meaning
$h_{c_{i}}^{*}$ Class representative embedding of node $i$ with label $c_{i}$, satisfying $\|C(h_{c_{i}}^{*})-c_{i}\|_{2}\leq\epsilon$
$C$ Trainable classifier in GNNs (e.g., MLP)
$\epsilon$ A small positive number used as a threshold and to ensure numerical stability
$c_{i}$ Class label of node $i$
$\hat{Y}$ Node predictions, $\hat{Y}=\text{GNN}(X,A)$
$X$ Node features
$A$ Adjacency matrix of the graph
GNN Graph Neural Network model, $\text{GNN}(X,A)=C(g(X,A))$
$g$ Aggregation function in GNNs (e.g., message-passing function)
$f$ Function to reconstruct node features based on existing predictions $\hat{Y}$
$Y$ Ground truth labels
$\hat{Y}_{t}$ Predictions at time $t$
$\hat{Y}_{t+1}$ Predictions at time $t+1$
$f_{\theta}$ Neural controller parameterized by $\theta$
$V$ Lyapunov function used to assess system stability
$V_{\phi}$ Neural Lyapunov function parameterized by $\phi$
$\Delta V$ Difference of the Lyapunov function between time steps, $\Delta V=V(C_{f_{\theta}}(\hat{Y}_{t}))-V(\hat{Y}_{t})$
$\Delta V_{\phi}$ Difference of the neural Lyapunov function, $\Delta V_{\phi}(\hat{Y}_{t})=V_{\phi}(C_{f_{\theta}}(\hat{Y}_{t}))-V_{\phi}(\hat{Y}_{t})$
$\mathcal{D}$ Domain of the control system
$\sigma$ ReLU activation function
$N$ Number of training samples
$\mathcal{L}_{\phi,\theta}$ Lyapunov loss function, $\mathcal{L}_{\phi,\theta}=\dfrac{1}{N}\sum\limits_{i=1}^{N}\left(\sigma(-V_{\phi}(\hat{Y}_{t}^{i}))+\sigma(\Delta V_{\phi}(\hat{Y}_{t}^{i}))\right)+V_{\phi}^{2}(Y)$
$r$ Replacement function used in test-time feature reconstruction
$\mathcal{C}$ Set of states violating Lyapunov conditions, $\mathcal{C}=\left\{\hat{Y}\ \bigg{|}\ \left(V_{\phi}(\hat{Y})\leq 0\ \cup\ \Delta V_{\phi}(\hat{Y})\geq 0\right)\cap\|\hat{Y}_{i}-Y_{i}\|_{2}\geq\epsilon\right\}$
$\theta$ Parameters of the neural controller $f_{\theta}$
$\phi$ Parameters of the neural Lyapunov function $V_{\phi}$
$C_{f_{\theta}}$ Composite function, $C_{f_{\theta}}(\hat{Y}_{t})=C(f_{\theta}(\hat{Y}_{t}))$
III Methodology
@@
In this section, we introduce the method of reconstructing node features using a neural controller. The overall process of this method is shown in Figure 1. Initially, we use a well-trained GNN to obtain the node predictions. This output is then fed back into the neural controller to reconstruct the node features, ensuring the new predictions remain close to the ground truth. Notably, the term ”node features” here refers explicitly to the original input features of the nodes. To ensure that the new predictions derived from the reconstructed node features remain stable near the ground truth, we introduce a neural Lyapunov function to aid in controller training. The variables and symbols used in this method are summarized in Table I. This section is organized as follows. We first explain the theoretical foundation of this method. Then, we detail the design of the neural controller and the neural Lyapunov function. Finally, we outline the training approach for this neural controller and the neural Lyapunov function.
III-A Problem Formulation
@@
In FRGNN  [10], the authors demonstrate that replacing the features of labeled nodes with class representative embeddings during the test phase can improve the model’s robustness to distribution shift. This paper provides the definition of class representative embeddings, as shown in Definition 1.
Definition 1.
@@
where $f$ is a function we aim to identify to minimize the objective function in Equation 3. However, directly solving this optimization problem poses significant challenges. The compositional nature of $f\circ g$ in GNNs introduces complex interactions between node features, graph topology, and the objective. This makes convergence to a global optimum (or even meaningful local optima) theoretically and practically difficult. Directly optimizing node features via gradient-based methods often leads to unstable dynamics, as small perturbations in feature space can propagate through message passing, causing erratic updates. To address this issue, we transform the optimization challenge into a control problem. We can formalize the graph neural network as a control system, treating node features as controlled variables and the predictions as states. However, merely formulating the problem as a control system is not sufficient to guarantee convergence to $Y$ or the stability of the system. This is where the concept of a Lyapunov function becomes essential. By introducing a Lyapunov function into our control design, we can provide a systematic way to analyze and ensure the stability of the system. Therefore, we propose to simultaneously optimize the control problem and learn an appropriate Lyapunov function. This joint optimization allows us to not only find a function $f$ that minimizes the objective in Equation 3 but also ensures that the system’s dynamics are stable and that the predictions converge to the ground truth $Y$. This approach leverages the strengths of control theory to overcome the limitations of directly solving the original optimization problem. The form of this control system is as follows:
@@
By transforming the optimization problem in Equation 3, we have shifted our focus to designing a controller $f$ and identifying a Lyapunov function. This approach ensures that the system output remains stable near the equilibrium point and that the predictions converge to the ground truth. In our framework, the equilibrium point is defined as the state where the system’s output remains unchanged, satisfying $C_{f_{\theta}}(Y)=Y$. When $Y$ is the ground truth, the system $C_{f_{\theta}}(\cdot)$ takes the true label data as input. The model’s predicted output will exactly match this real label, resulting in no prediction error that needs further correction. Consequently, the system’s next output $C_{f_{\theta}}(Y)$ equals the current input $Y$, thus satisfying the equilibrium condition. By defining equilibrium in this way, we can leverage stability analysis to ensure that the system will drive the predictions toward the ground truth over time. In the following sections, we detail the design of the neural controller and the neural Lyapunov function and introduce the training methodology for these components.
III-B Theoretical Foundation
In the previous section, we transform the GNNs optimization problem into a control problem. Now, the main challenge is designing a controller $f$ to adjust the node features so that the new predictions derived from these reconstructed features get close to the ground truth. Traditional controllers, such as PID, LQR, and AC, are not well-suited for this system due to the complex nonlinearity of graph neural networks. Moreover, unlike ordinary neural networks like MLPs, GNNs include message-passing layers, meaning that modifying a specific node feature not only affects the prediction of that node but also influences the predictions of other nodes through the graph structure. This coupling makes the control problem particularly challenging for traditional controllers. Therefore, we choose a neural controller, leveraging the powerful representation capabilities of neural networks to learn an effective controller suitable for this system. We denote the neural controller as $f_{\theta}$, where $\theta$ represents the trainable parameters. Thus, Equation 5 can be rewritten in closed-loop form as follows:
@@ III-B Theoretical Foundation
Theorem 1.
Consider a closed-loop system $\hat{Y}_{t+1}=C_{f_{\theta}}(\hat{Y_{t}})$, where $C_{f_{\theta}}$ is a MLP with ReLU activation function. Consider the system has an equilibrium point at the ground truth $Y$. Assume there exists a continuously differentiable function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies the following conditions:
@@
Proof.
Please refer to the Appendix A. ∎
Remark (Lyapunov Function and Stability Criterion) Consider a continuously differentiable function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that maps the state space to a scalar value. For a closed-loop system $\hat{Y}_{t+1}=C_{f}^{\theta}(\hat{Y}_{t})$ with an equilibrium point at $Y^{*}$, the system is asymptotically stable at $Y^{*}$ if $V$ satisfies: (1) $V(Y^{*})=0$, (2) $V(\hat{Y})>0$ for all $\hat{Y}\neq Y^{*}$, and (3) $\Delta V(\hat{Y})=V(\hat{Y}_{t+1})-V(\hat{Y}_{t})<0$ for all $\hat{Y}\neq Y^{*}$. Here $V$ is called a Lyapunov function for the system.
@@
Figure 2: The architecture of the neural controller and the neural Lyapunov function and Lyapunov loss. We simultaneously train the neural Lyapunov function and the neural controller to obtain a control system that ensures stability and effectively guides the GNN predictions to gradually approach the ground truth.
III-C Training Methodology